# Deep Bayesian Active Learning for Preference Modeling in Large Language Models

### v1

@@ 1 Introduction
Preference Modeling is a key component to aligning unsupervised pre-trained Large Language Models (LLMs) towards human preferences [51, 42, 2, 54]. It is often performed by collecting human feedback for a set of prompt-answer pairs and then leveraging the data to steer the behavior of such models, either directly [45] or via reward models [59]. Nevertheless, human feedback generation is laborious [7], especially when it requires specialized knowledge [3, 22]. Furthermore, the quality of the prompts has a crucial impact on the performance of fine-tuned models [52]. Hence, selecting the most informative points to gather feedback is essential to reduce costs and enable better LLMs.
Despite its substantial impact, data selection for Preference Modeling poses a significant challenge. The prompt-answer pool is arbitrarily large and semantically rich. Additionally, human feedback is inherently noisy, with low agreement rates among labelers, typically between 60% – 75% [59, 51, 12, 10]. Lastly, the intrinsic scale of LLM development requires parallelized labeling and makes frequent model updates prohibitively expensive, limiting the applicability of many active learning schemes that rely on single-point acquisition [28].
@@
Figure 2: An illustration of how BAL-PM works. For each tuple $(x,y_{1},y_{2})\in\mathcal{D}_{pool}$, we obtain features for the prompt and prompt-answer pairs by computing the last layer embeddings of the base LLM. We leverage the prompt feature space to estimate the entropy score of the acquired prompt distribution, ^$\hat{\mathcal{H}}({X}_{train}\cup\{x\})$. Similarly, we use the prompt-answer features as input for the Bayesian Preference Model, which is used to estimate task-dependent epistemic uncertainty scores, $\hat{U}(x,y_{1},y_{2})$. BAL-PM selects the tuple that maximizes the linear combination of both scores.
@@ 2 Related Work
Active Preference Modeling aims to leverage active learning techniques and reduces the volume of feedback needed for Preference Modeling [23]. There has been a recent surge of interest in the area [39, 25, 1, 11, 18, 13] given the impact of Preference Optimization in fine-tuning LLMs [42, 45, 52, 59]. A portion of this line of work focuses on query generation to directly optimize preferences [39, 25, 11, 13]. While having the same goal of reducing the human feedback needed for learning, this configuration is orthogonal to ours. We focus on the pool-based setting, as in Gleave and Irving [18], which allows us to leverage real human feedback in experiments rather than relying on synthetic preference simulators. Gleave and Irving [18] was the first attempt of Bayesian Active Learning in this setting, and it directly applies BALD acquisition [35] for Preference Modeling. In contrast, our work proposes a new objective that extends BALD acquisition to account for the entropy of the acquired prompt distribution to encourage the acquisition of more diversified samples.
@@ 3 Preliminaries
Problem Statement. We formulate our setup as an active inverse variant of the contextual dueling bandit problem [49, 46]. We assume a prompt space $\mathcal{X}$, an action space $\mathcal{Y}$, and a language policy $\tau:\mathcal{X}\times\mathcal{Y}\rightarrow[0,\infty)$. Given $x\sim\mathcal{X}$, this language policy $\tau$ (for instance, a LLM) selects actions $y_{1},y_{2}\sim\tau(\cdot\mid x)$ (also referred to as answers), generating a dataset of tuples $\mathcal{D}_{pool}=\{x^{i},y_{1}^{i},y_{2}^{i}\}^{N}$. Crucially, $x$ is sampled with replacement, i.e., we may generate multiple completions for the same prompt. Then, we define a policy $\pi:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$, which select tuples $(x,y_{1},y_{2})\in\mathcal{D}_{pool}$ to query for human binary preference over completions $y_{1}\succ y_{2}$, forming a preference dataset $\mathcal{D}_{train}=\{x^{i},y_{1}^{i},y_{2}^{i},{y_{1}\succ y_{2}}^{i}\}^{B}$. Finally, $D_{train}$ is used to learn a preference model $p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})$, parameterized by $\boldsymbol{\theta}$, which aims to recover the human preference function. The goal is to find $\pi$ that minimizes the amount of samples $B$ required to learn $p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})$.
@@ 4 Bayesian Active Learner for Preference Modeling
We now introduce our method for Active Preference Modeling, BAL-PM, illustrated in Figure 2. Our desiderata is to design an acquisition policy that addresses the shortcomings of naive epistemic uncertainty estimation – such as the acquisition of redundant samples – by leveraging a unsupervised source of epistemic uncertainty that encourages diversity in the acquired training distribution.
@@
However, naively applying the KL estimator from Equation 4 has a major drawback: the training set $\mathcal{D}_{train}$ initially contains very few data points and does not provide enough support to represent the probability density, introducing bias to the estimates and affecting the data selection.
For illustration, we show the scenario of Figure 3 (a). In this case, we estimate the entropy using Equation 4, with $k$ = 3. Since it does not account for the available points in the pool, it underestimates the density around the top cluster and ends up selecting the green point as the one that maximizes the entropy of the feature space, while the point that does so is in the bottom cluster. In an extreme case where all the points in the top cluster are the same, this bias leads to acquiring duplicated points. In Appendix E we formally derive the KL entropy estimator and show how the low-data regime challenges its main assumptions.
@@
Equation 11 acquire points by computing $D(x)$ (based on the kNN distance) and the counter $n_{X_{tr}}$ related to prompt $x$ only. Furthermore, as $D(x)$ accounts for the full dataset in the KSG estimator, it does not change over training. Hence, we may compute it offline once, and potentially scale to very large datasets [8]. Lastly, BAL-PM acquisition scheme builds a batch of data by successively selecting points following Equation 5. Crucially, while BAL-PM keeps the preference model uncertainty estimates the same over the batch, it updates the entropy term after in-batch iteration. This operation boils down to updating the counter $n_{X_{tr}}$, a lightweight operation. In 1, we present the pseudocode for BAL-PM.
@@ 5 Experiments and Discussion
Figure 4: Comparison with baseline methods in Active Preference Modeling. BAL-PM considerably reduces the number of samples required for preference modeling, achieving 33% and 68% of reduction in the Reddit TL;DR test split and CNN/DM News datasets, respectively.
Experimental Setup. We consider a pool-based active learning setup. Each experiment consists of several acquisition cycles, where each iteration performs a batch acquisition in the currently available pool. The training set starts with the size of one acquired batch and leaves the remaining data for the pool set. Following previous works [28, 16], we reinitialize the model after each acquisition to decorrelate subsequent acquisitions. We train the ensemble of adapters on previously acquired data and employ early stopping based on the log-likelihood of a held-out validation set. We evaluate the preference model after each acquisition loop and report the average log-likelihood of the test sets. Finally, Appendix C details all hyperparameters and tuning methodology used in this work. [^2] We release our code at https://github.com/luckeciano/BAL-PM.
@@
Datasets. Following previous work [18], we considered prompts from the Reddit TL;DR dataset of Reddit posts [56] and the CNN/DM News dataset [21]. We leverage the generated answers and human feedback collected by Stiennon et al. [51]. The Reddit dataset contains train/eval/test splits, and we adopt the train split (92,858 points) for the pool and training sets, the eval split (33,083 points) for validation, and report results in the test set (50,719 points). The CNN/DM dataset contains a single split (2,284 points), and we use it for the Out-Of-Distribution (OOD) evaluation.
@@ 5.1 Experiments
The middle plot shows the ratio of unique prompts in each active learning loop, and BAL-PM acquires batches with all distinct prompts during almost the whole training. BALD only maintains a rate of 70%, which means a substantial number of duplicated prompts. In Appendix G, we present the first batch sampled by BALD and BAL-PM for a qualitative analysis. Lastly, the plot on the right shows the ratio of unique prompts across all training. While random sampling presents a high unique prompt ratio in each separate batch, it consistently samples duplicated prompts throughout learning. In contrast, BAL-PM maintains a high ratio of unique prompts during most of the training. Again, this rate progressively decays as BAL-PM exhausts the pool of different prompts and due to the influence of the epistemic uncertainty prioritizing particular prompt-answer pairs.
@@
Ablations and Further Analysis. We conduct ablation studies in the key components of the proposed method in Appendix D. More concretely, we ablate the components of the objective to show that both preference model epistemic uncertainty and entropy scores play a relevant role in BAL-PM. We also ablate the type of uncertainty and the employed entropy estimator. Lastly, we conduct further empirical analysis in Appendix F to investigate how each component of Equation 5 contributes to the data selection.
@@ 6 Closing Remarks
We thank Jannik Kossen for the insightful discussions in the early stages of this project. Luckeciano C. Melo acknowledges funding from the Air Force Office of Scientific Research (AFOSR), award number FA8655-21-1-7017.
@@ Appendix F BAL-PM Objective – Empirical Analysis
Appendix G Batch Samples

### v2

@@ 1 Introduction
Preference Modeling is a key component to aligning unsupervised pre-trained Large Language Models (LLMs) towards human preferences [1, 2, 3, 4]. It is often performed by collecting human feedback for a set of prompt-completion pairs and then leveraging the data to steer the behavior of such models, either directly [5] or via reward models [6]. Nevertheless, human feedback generation is laborious [7], especially when it requires specialized knowledge [8, 9]. Furthermore, the quality of the prompts has a crucial impact on the performance of fine-tuned models [10]. Hence, selecting the most informative points to gather feedback is essential to reduce costs and enable better LLMs.
Despite its substantial impact, data selection for Preference Modeling poses a significant challenge. The prompt-completion pool is arbitrarily large and semantically rich. Additionally, human feedback is inherently noisy, with low agreement rates among labelers, typically observed between 60% – 75% for these settings [6, 1, 11, 12]. Lastly, the intrinsic scale of LLM development requires parallelized labeling and makes frequent model updates prohibitively expensive, limiting the applicability of many active learning schemes that rely on single-point acquisition [13].
@@
Figure 2: An illustration of how BAL-PM works. For each tuple $(x,y_{1},y_{2})\in\mathcal{D}_{pool}$, we obtain features for the prompt and prompt-completion pairs by computing the last layer embeddings of the base LLM. We leverage the prompt feature space to estimate the entropy score of the acquired prompt distribution, ^$\hat{\mathcal{H}}({X}_{train}\cup\{x\})$. Similarly, we use the prompt-completion features as input for the Bayesian Preference Model, which is used to estimate task-dependent epistemic uncertainty scores, $\hat{U}(x,y_{1},y_{2})$. BAL-PM selects the tuple that maximizes the linear combination of both scores.
@@ 2 Related Work
Active Preference Modeling leverages active learning techniques to reduce the feedback needed for Preference Modeling [27]. There has been a recent surge of interest in the area [32, 33, 34, 35, 20, 36] given the impact of Preference Optimization in fine-tuning LLMs [2, 5, 10, 6]. A portion of this line of work focuses on query generation to directly optimize preferences. Mehta et al. [32] theoretically formalizes the problem and proposes a method that generates one completion to maximize the uncertainty of the triple in a kernelized setting. Das et al. [35] proposes a method based on confidence bands, accounting for both completions in the triple and relaxing linearity assumptions on the reward function. Ji et al. [33] constructs an optimistic estimator for the reward gap between completions and selects those with the least gap, using an uncertainty estimator to reduce query complexity. Lastly, Dwaracherla et al. [36] generate completions using double Thompson Sampling, representing epistemic uncertainty with an Epistemic Neural Network [37], similar to our Bayesian model. Overall, while having the shared goal of reducing the volume of human feedback, query generation is orthogonal to our problem setting. Instead, we focus on the pool-based setting, as in Gleave and Irving [20], which allows us to leverage real human feedback in experiments rather than relying on synthetic preference simulators. Gleave and Irving [20] was the first attempt of Bayesian Active Learning in this setting, and it directly applies BALD acquisition [23] for Preference Modeling, using a fully fine-tuned deep ensemble for epistemic uncertainty estimation. In contrast, our work proposes a new objective that extends BALD acquisition to account for the entropy of the acquired prompt distribution to encourage the acquisition of more diversified samples, and formulates the Bayesian model as an ensemble of adapters.
@@ 3 Preliminaries
Problem Statement. We formulate our setup as an active inverse variant of the contextual dueling bandit problem [45, 46]. We assume a prompt space $\mathcal{X}$, an action space $\mathcal{Y}$, and a language policy $\tau:\mathcal{X}\times\mathcal{Y}\rightarrow[0,\infty)$. Given $x\sim\mathcal{X}$, this language policy $\tau$ (e.g., an LLM) selects actions $y_{1},y_{2}\sim\tau(\cdot\mid x)$ (also referred to as completions), generating a dataset of tuples $\mathcal{D}_{pool}=\{x^{i},y_{1}^{i},y_{2}^{i}\}^{N}$. Crucially, $x$ is sampled with replacement, i.e., we may generate multiple completions for the same prompt. Then, we define a policy $\pi:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow[0,\infty)$, which select tuples $(x,y_{1},y_{2})\in\mathcal{D}_{pool}$ to query for human binary preference over completions $y_{1}\succ y_{2}$, forming a preference dataset $\mathcal{D}_{train}=\{x^{i},y_{1}^{i},y_{2}^{i},{y_{1}\succ y_{2}}^{i}\}^{B}$. Finally, $D_{train}$ is used to learn a preference model $p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})$, parameterized by $\boldsymbol{\theta}$, which aims to recover the human preference function. The goal is to find $\pi$ that minimizes the amount of samples $B$ required to learn $p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})$.
@@ 4 Bayesian Active Learner for Preference Modeling
We now introduce our method for Active Preference Modeling, BAL-PM, illustrated in Figure 2. Our desiderata is to design an acquisition policy that addresses the shortcomings of naive epistemic uncertainty estimation – such as the acquisition of redundant samples – by leveraging an unsupervised source of epistemic uncertainty that encourages diversity in the acquired training distribution.
@@
However, naively applying the KL estimator from Equation 4 has a major drawback: the training set $\mathcal{D}_{train}$ initially contains very few data points and does not provide support to represent the probability density, introducing bias to the estimates and affecting the data selection.
For illustration, we show the scenario of Figure 3a. In this case, we estimate the entropy using Equation 4, with $k$ = 3. Since it does not account for the available points in the pool, it underestimates the density around the top cluster and ends up selecting the green point as the one that maximizes the entropy of the feature space, while the point that does so is in the bottom cluster. In an extreme case where all the points in the top cluster are the same, this bias leads to acquiring duplicated points. In Appendix E we formally derive the KL entropy estimator and show how the low-data regime challenges its main assumptions.
@@
Equation 11 acquire points by computing $D(x)$ (based on the kNN distance) and the counter $n_{X_{tr}}$ related to prompt $x$ only. Furthermore, as $D(x)$ accounts for the full dataset in the KSG estimator, it does not change over training. Hence, we may compute it offline once, and potentially scale to very large datasets [55]. Lastly, BAL-PM acquisition scheme builds a batch of data by successively selecting points following Equation 5. Crucially, while BAL-PM keeps the preference model uncertainty estimates the same over the batch, it updates the entropy term after in-batch iteration. This operation boils down to updating the counter $n_{X_{tr}}$, a lightweight operation. In Algorithm 1, we present the pseudocode for BAL-PM. Appendix H further describes its computational cost.
@@ 5 Experiments and Discussion
Figure 4: Comparison with baseline methods in Active Preference Modeling. BAL-PM considerably reduces the number of samples required for preference modeling, achieving 33% and 68% of reduction in the Reddit TL;DR test split and CNN/DM News datasets, respectively. The shaded area corresponds to the standard error computed over five seeds.
Experimental Setup. We consider a pool-based active learning setup. Each experiment consists of several acquisition cycles, where each iteration performs a batch acquisition in the currently available pool. The training set starts with the size of one acquired batch and leaves the remaining data for the pool set. Following previous works [13, 14], we reinitialize the model after each acquisition to decorrelate subsequent acquisitions. We train the ensemble of adapters on previously acquired data and employ early stopping based on the Log-Likelihood of a held-out validation set. We evaluate the preference model after each acquisition loop and report the average Log-Likelihood of the test sets. Appendix G discusses why the test average Log-Likelihood is a proper metric for Preference Modeling. Finally, Appendix C details hyperparameters and tuning methodology used in this work [^2] We release our code at https://github.com/luckeciano/BAL-PM..
@@
Datasets. Following previous work [20], we considered prompts from the Reddit TL;DR dataset of Reddit posts [25] and the CNN/DM News dataset [26]. We leverage the generated completions and human feedback collected by Stiennon et al. [1]. The Reddit dataset contains train/eval/test splits, and we adopt the train split (92,858 points) for the pool and training sets, the eval split (33,083 points) for validation, and report results in the test set (50,719 points). The CNN/DM dataset contains a single split (2,284 points), and we use it for the Out-Of-Distribution (OOD) evaluation.
@@ 5.1 Experiments
The middle plot shows the ratio of unique prompts in each active learning loop, and BAL-PM acquires batches with all distinct prompts during almost the whole training. BALD only maintains a rate of 70%, which means a substantial number of duplicated prompts. In Appendix K, we present the first batch sampled by BALD and BAL-PM for a qualitative analysis. Lastly, the plot on the right shows the ratio of unique prompts across all training. While random sampling presents a high unique prompt ratio in each separate batch, it consistently samples duplicated prompts throughout learning. In contrast, BAL-PM maintains a high ratio of unique prompts during most of the training. Again, this rate progressively decays as BAL-PM exhausts the pool of different prompts and due to the influence of the epistemic uncertainty prioritizing particular prompt-completion pairs.
@@
Ablations and Further Analysis. We conduct ablation studies in the key components of the proposed method in Appendix D. More concretely, we ablate the components of the objective to show that both preference model epistemic uncertainty and entropy scores play a relevant role in BAL-PM. We also ablate the type of uncertainty and the employed entropy estimator. Furthermore, we conduct further empirical analysis in Appendix F to investigate how each component of Equation 5 contributes to the data selection, and conduct a robustness analysis for the $\beta$ hyperparameter in Appendix I. Lastly, we provide comparison with additional data selection baselines in Appendix J.
@@ 6 Closing Remarks
We thank Jannik Kossen for the insightful discussions in the early stages of this project. We also thank the reviewers for providing insightful feedback. Luckeciano C. Melo acknowledges funding from the Air Force Office of Scientific Research (AFOSR) European Office of Aerospace Research & Development (EOARD) under grant number FA8655-21-1-7017.
@@ Appendix F BAL-PM Objective – Empirical Analysis
Appendix G Is Log-Likelihood a proper performance measure for Preference Modeling?
In this Section, we argue why the Average Log-Likelihood on the test set is a good performance measure for Preference Modeling. Given a test set $\mathcal{D}_{test}=\{(x,y_{1},y_{2},y_{1}\succ y_{2})\}^{N}$ and the learned preference model $p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})$, the average Log-Likelihood is given by:
$LL(\mathcal{D}_{test},\boldsymbol{\theta})=\mathbb{E}_{(x,y_{1},y_{2},y_{1}\succ y_{2})\sim\mathcal{D}_{test}}[\log{p_{\boldsymbol{\theta}}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})}].$ (17)
Equation 17 is exactly the objective maximized in standard binary classification (or, equally, the minimization of the negative Log-Likelihood loss) but computed over the test data. In other words, this is the negative "test loss".
Average LL is a typical metric in the Active Learning and Uncertainty Quantification literature [61, 62, 63]. For Preference Modeling, it is very relevant as LL directly accounts for the preference strength to rank models: given a triple $(x,y_{1},y_{2})$ where all raters agree that $y_{1}$ is preferable over $y_{2}$, LL allows us to measure that a model A predicting $p_{A}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})=0.9$, ($LL\approx-0.1)$ is better (in that data point) than another model B predicting $p_{B}(y_{1}\succ y_{2}\mid x,y_{1},y_{2})=0.6$ ($LL\approx-0.5)$. Accuracy would provide an equal score for both models since it only accounts for the binarized prediction. LL provides a more "fine-grained" measure.
Another crucial point is that LL factors in the aleatoric uncertainty in the label-generating process. For instance, in a scenario where only 70% of the raters agree that $y_{1}$ is preferable, LL better ranks models whose predictions are closer to $p=0.7$, respecting the ground truth preference strength, which is not possible with accuracy.
G.1 Do the models better ranked by Average Log Likelihood (LL) lead to better fine-tuned policies?
In the context of Preference Modeling, fine-tuning LM policies is currently a very relevant downstream task. The Preference Modeling optimization objective and model selection protocol adopted in this work follow exactly the prior influential work on the topic [6, 1], which provides evidence that better preference models (in terms of validation loss) lead to improved downstream policies. Thus, we expect our models to behave similarly under the same conditions.
Figure 12: The Relationship between the Test Log-Likelihood of a Preference Model and the Performance of the corresponding fine-tuned Policy. We show that, under simple conditions, there is a strong correlation between these two performance measures.
As additional evidence, we empirically illustrate the relationship between Log-Likelihood and policy performance on a simplified setup (Figure 12). Here, prompts $x$ and completions $y$ are real numbers in $[0,1]$. The ground-truth reward function is given by a Gaussian density function $r(x,y)=\mathcal{N}(x+y\mid\mu=1.0,\sigma=0.4)$, and true preferences follow the Bradley-Terry model. In this setup, we progressively increase the training set size (the x-axis in Figure 12(a)) at which we train the preference models. This process generates different models with increasing levels of test-set average Log-Likelihood. Then, similar to [64], we optimize the base policy via a Best-of-N optimizer by leveraging each of these learned preference models. Finally, we report the rate at which the fine-tuned policy completion is preferable over the base policy completion according to the ground truth reward model ("win rate"). Although simple, this setting allows us to bypass several optimization and distributional challenges and solely focus on evaluating the relationship between average Log-Likelihood and the performance of the fine-tuned policy. Figure 12(a) reports the Log-Likelihood (red) and the win rate against the base policy (blue). Figure 12(b) directly plots both measures and fits a regression line. We observe a strong correlation, which aligns with our point: a higher test-set average Log-Likelihood means that the preference model is better at predicting the ground truth preferences, assigning higher rewards for better completions, and, therefore, improving fine-tuned policies that maximize such reward scores.
Appendix H BAL-PM Computational Cost Description
In this section, we describe the computational cost of executing BAL-PM. We argue that computational tractability is one of the main contributions of our method. We start by providing some context: our work focuses on (Bayesian) Active Learning, which is naturally more computationally demanding than simply training predictive models. This is because we require models that express epistemic uncertainty to acquire informative labels for efficient training. This also requires models to constantly update their uncertainties based on the new data via re-training. The key is that Active Learning reduces the number of labels needed to train a better model, which considerably overcomes the extra computational cost. Labeling is significantly more expensive and laborious.
As described in Section 1, Preference Modeling in LLMs requires batch acquisition – it is impossible to request the label of a single point, re-train the model, and repeat this process. Still, tractable methods rely on these single-point acquisition objectives. Thus, what BAL-PM does computationally is to replace $B$ model re-trainings per acquired batch with computing entropy estimates (considerably cheaper, as explained below). $B$ is the batch size, and we set $B=320$ in our experiments.
BAL-PM does not require training or inference on LLMs during the active learning loops. This considerably reduces the computational cost and allows us to scale up to 140b models in a single A100 GPU. Comparatively, even fully fine-tuning a 7b-parameter model currently requires at least several A100 GPUs. LoRA methods [65] also require new LLM inferences for every model update, while BAL-PM only requires a single time.
The computation of BAL-PM has three pieces: offline processing (LLM inference and kNN computation), updating adapters, and entropy estimation. LLM inference is done only once before Active Learning, which is the minimum for LLM adoption. Furthermore, we may compute the features used for the preference model and sentropy estimation in the same forward pass: every prompt-completion input concatenates prompt/completion texts. Thus, we can extract prompt features as the last layer embedding right after the last prompt token and the prompt-completion features right after the completion’s last token. Hence, there is no extra cost to extract features for entropy estimation. The cost of updating adapters is minimal: it boils down to updating MLPs with two hidden layers, which is reasonably cheap for LLM research. Lastly, the entropy estimation only requires computing the di-gamma function (Equation 11) in the pool.
Appendix I $\beta$ Robustness Analysis
In this Section, we introduce a robustness analysis for the $\beta$ hyperparameter (Figure 8) considering the values in the search space, described in Table 2. As presented in Equation 5, this hyperparameter balances the effect of the epistemic uncertainty and entropy scores.
(a) Reddit TL;DR (Test)
(b) CNN/DM Dataset (OOD)
Figure 13: $\beta$ Robustness Analysis. We considered different scales of the $\beta$ hyperparameter in the BAL-PM objective (Equation 5).
The impact of the choice is more noticeable when values are 100x greater/lower than the optimal choice. Values around 10x greater/lower still perform well, suggesting good room for choosing this hyperparameter. Furthermore, we employed the same value of $\beta$ across the different datasets and LLMs, suggesting robustness across different relevant dimensions. Crucially, $\beta$ trades off the contribution of two different terms. As such, it provides a spectrum of objectives and may recover the two extremes presented in the ablation of Figure 13. Naturally, different choices of $\beta$ will change the uncertainty score ratio in Figure 11 on Appendix F (i.e., the contribution of each term after convergence). Nevertheless, and most importantly, the behavior of the curves –- the entropy contribution progressively reducing and converging and the relevance of epistemic uncertainty estimates increasing –- should remain.
Appendix J Further Baselines
Figure 14: Comparison with Preference Model trained on the full dataset.
In this section, we provide additional baselines for further comparison.
J.1 Full Dataset Baseline
We start by evaluating the performance of a preference model trained in the full dataset. Figure 14 presents this result in purple, with the shaded area representing the standard error computed across five seeds. BAL-PM achieves on-par performance to this baseline, although it only requires  24000 data points (the full dataset contains 92,858 points). This result is another strong evidence of the sample efficiency of our method.
J.2 Additional Data Selection Methods
(a) Reddit TL;DR (Test)
(b) CNN/DM Dataset (OOD)
Figure 15: Comparison with several additional baselines for Active Preference Modeling. These baselines focus on different notions of uncertainty and diversity for acquiring samples.
We considered the following additional baselines:
  • Entropy Minimizer: Entropy Minimizer: Inspired by Liu et al. [66], we consider an objective that, in addition to selecting points with high epistemic uncertainty, also selects points that are semantically similar to the current training points. This is equivalent to selecting points that increase the entropy of the prompt distribution the least, thus the name "Entropy Minimizer". It serves as a check for our central hypothesis that entropy maximization leads to better batch active learning.
  • Perplexity: Inspired by Gonen et al. [67], we consider an objective that selects points based on the perplexity of the base LLM. We consider two versions: one that chooses points with lower perplexity (Low Perplexity) and another with higher perplexity (High Perplexity). This is an interesting baseline since perplexity is equivalent to the predictive entropy of the token distribution. Therefore, it helps to analyze how much the base LLM "knows what it does not know" in terms of preference modeling.
  • MC Dropout [63]: This method performs approximate Bayesian inference via executing dropout at test time to generate different parameter hypotheses. Therefore, it can express epistemic uncertainty, which is used to select the most informative points.
  • Latent Reward Uncertainty (LRU): This method computes a reward distribution over the data points by leveraging the latent reward model learned via the Bradley-Terry model. Then, it selects extreme points (too high or too low rewards) as a proxy for the uncertainty of the model.
(a) Reddit TL;DR (Test)
(b) CNN/DM Dataset (OOD)
Figure 16: The effect of incorporating the entropy objective in uncertainty baselines. This shows how our proposed objective can also boost the performance of different baselines.
Figure 15 reports performances for both test and OOD sets. In both cases, BAL-PM outperforms additional baselines. In the sequence, MC-Dropout works best as the baseline that targets the epistemic uncertainty of a Bayesian model. Unsurprisingly, Entropy Minimizer and Low Perplexity work worse since they target points with lower entropy. LRU presented mixed results, suggesting that the latent reward may not represent well the preference model’s uncertainty. More interestingly, while these models can represent different uncertainties to seek informative points, they naturally cannot provide in-batch diversity - they suffer from the same challenges as BALD. In this perspective, the BAL-PM objective can also improve upon those methods, as we show in Figure 16: we combined MC-Dropout and LRU with our entropy term to provide in-batch diversity, which consistently improved both methods across the datasets.
Appendix K Batch Samples